Find 3 consecutive integers so that 3 times the first increased by twice the
second is 2 less than 4 times the third one.
My idea starts: 3x + 2x = 4x - 2

Answer by sachi(548)

let the 3 consecutive integers are x-1 ,x & x+1
now 3 times the first increased by twice the second is 2 less than 4 times the third one
so 3(x-1)+2x=4(x+1)-2
or 3x-3+2x=4x+4-2
or x=5
so the 3 consecutive integers are 4,5,6
ans